# Finite-time Scaling beyond the Kibble-Zurek Prerequisite: Driven Critical Dynamics in Strongly Interacting Dirac Systems

### v1

@@ I IMAGINARY-TIME DRIVEN DYNAMICS
The driven dynamics is realized by varying $U$ with imaginary time $\tau$ as $U=U_{0}+R\tau$, starting from the ground state at $U_{0}$. The wave function $|\psi(\tau)\rangle$ obeys the imaginary-time Schrödinger equation De Grandi et al. (2011); Grandi et al. (2013); Liu et al. (2013)
@@ II Determinant quantum Monte Carlo
We employ the large-scale determinant QMC method Assaad and Evertz (2008) to investigate the imaginary-time driven dynamics of our models. Specifically, we start with a trial wave function |$\ket{\psi_{T}}$ , and project the system onto the desired initial state |$\ket{\psi_{0}}$ with a sufficiently long projection time $\tau_{0}$. Then we vary the system parameter linearly with imaginary time variable $\tau$ and observe the scaling behavior of observables at final state, at which the expectation value of observables is given by
$\braket{O(\tau)}=\frac{\bra{\psi_{0}}U(2\tau,\tau)~{}O~{}U(\tau,0)\ket{\psi_{0}}}{\bra{\psi_{0}}U(2\tau,0)\ket{\psi_{0}}}=\frac{\bra{\psi_{T}}\mathrm{e}^{-\tau_{0}H_{\mathrm{initial}}}U(2\tau,\tau)~{}O~{}U(\tau,0)\mathrm{e}^{-\tau_{0}H_{\mathrm{initial}}}\ket{\psi_{T}}}{\bra{\psi_{T}}\mathrm{e}^{-\tau_{0}H_{\mathrm{initial}}}U(2\tau,0)\mathrm{e}^{-\tau_{0}H_{\mathrm{initial}}}\ket{\psi_{T}}}.$ (S2)
In numerical calculations, we use Trotter decomposition to discretize imaginary-time propagator into ${M=\Theta/\Delta\tau}$ (${M}$ is integer and $\Theta$ is imaginary time) time slices with
$\mathrm{e}^{-\Theta H}=\prod^{M}_{m=1}\left[\mathrm{e}^{-\Delta\tau H_{t}}\mathrm{e}^{-\Delta\tau H_{U}}+\mathcal{O}\left(\Delta\tau^{2}\right)\right],$ (S3)
where $H_{t}$ and $H_{U}$ are the hopping term and the interaction term respectively in the Hamiltonian. We choose small enough $\Delta\tau/t\leq 0.05$. To decouple two-body fermion-fermion coupling form of $\mathrm{e}^{\Delta\tau H_{U}}$, we use discrete Hubbard-Stratonovich transformation
@@
$\mathrm{e}^{-\frac{\Delta\tau U}{2}\left(n_{i\uparrow}+n_{i\downarrow}-1\right)^{2}}=\sum_{l=\pm 1,\pm 2}\gamma(l)\mathrm{e}^{\mathrm{i}\sqrt{\frac{\Delta\tau U}{2}}\eta(l)\left(n_{i\uparrow}+n_{i\downarrow}-1\right)},$ (S4)
$\mathrm{e}^{-\frac{\Delta\tau V}{2}\left(n_{i}+n_{j}-1\right)^{2}}=\sum_{l=\pm 1,\pm 2}\gamma(l)\mathrm{e}^{\mathrm{i}\sqrt{\frac{\Delta\tau V}{2}}\eta(l)\left(n_{i}+n_{j}-1\right)},$ (S5)
to obtain one-body fermion-auxiliary field coupling. Here, we introduce four-component space-time local auxiliary fields $\gamma(\pm 1)=1+\sqrt{6}/3$, $\gamma(\pm 2)=1-\sqrt{6}/3$, $\eta(\pm 1)=\pm\sqrt{2\left(3-\sqrt{6}\right)}$, $\eta(\pm 2)=\pm\sqrt{2\left(3+\sqrt{6}\right)}$, and use DQMC for importance sampling over these space-time configurations. Next, we elaborate on how DQMC numerically calculates the sampling weight.
Taking Hubbard interaction as an example, for each imaginary time and each position of the interaction, we employ Hubbard-Stratonovich transformation as in Eq. (S4). This means that we introduce an auxiliary field in $d+1$ dimensions. As a result, the imaginary-time propagator can be fully expressed using single-particle operators. This allows us to represent it in the following quadratic form of fermion operators:
$\mathrm{e}^{-\Theta H}\equiv\sum_{\mathrm{c}}\mathrm{e}^{-\Theta H_{\mathrm{c}}}=\sum_{\mathrm{c}}A_{\mathrm{c}}\prod^{M}_{m=1}~{}\mathrm{e}^{\vec{c}^{\dagger}T\vec{c}}~{}\mathrm{e}^{\vec{c}^{\dagger}V_{\mathrm{c}(m)}\vec{c}},$ (S6)
where $\sum_{\mathrm{c}}$ denotes the summation over all space-time configurations of the auxiliary field. Considering that each local component of the auxiliary field has 4 possible values, the summation comprises up to $4^{MN}$ terms, where $N$ represents the number of spatial degrees of freedom. $H_{c}$ denotes the decoupled configuration Hamiltonian, while $T$ and $V_{\mathrm{c}(m)}$ are the resulting quadratic coefficient matrices from the rearrangement, and $A_{\mathrm{c}}$ is the coefficient. Both $V_{\mathrm{c}(m)}$ and $A_{\mathrm{c}}$ depend on the auxiliary field configuration. The complete form of the evolution operator has been presented above. Next, we consider expressing the trial wave function. We use a DSM inital state for trial wave function, which is a direct product state, and numerically it can be written as the following Slater determinant:
|$\ket{\psi_{T}}=\bigotimes^{N_{\mathrm{e}}}_{n_{\mathrm{e}}=1}\left[\left(\sum_{x}c_{x}^{\dagger}P_{x,n_{\mathrm{e}}}\right)\ket{0}\right]=\bigotimes^{N_{\mathrm{e}}}_{n_{\mathrm{e}}=1}\left[\left(\vec{c}^{\dagger}P\right)_{n_{\mathrm{e}}}\ket{0}\right],$ (S7)
where $N_{\mathrm{e}}$ denotes the number of electrons. This implies that the trial wave function is a direct product of $N_{\mathrm{e}}$ fermion single-particle wave functions. The index $x$ denotes the degree of freedom of the electron, including spatial degrees of freedom, spin degrees of freedom, etc. The matrix element $P_{x,n_{\mathrm{e}}}$ represents the probability amplitude of the $n_{\mathrm{e}}$th electron on the $x$th degree of freedom. Note that the imaginary-time propagator $\mathrm{e}^{-\Theta H_{\mathrm{c}}}$ is essentially the Boltzmann weight factor of the auxiliary field configuration in statistical mechanics. According to Eqs. (S6) and (S7), the partition function of the auxiliary field configuration can be expressed as:
$Z=\sum_{\mathrm{c}}\bra{\psi_{T}}\mathrm{e}^{-\Theta H_{\mathrm{c}}}\ket{\psi_{T}}=\sum_{\mathrm{c}}A_{\mathrm{c}}~{}\mathrm{det}\left[P^{\dagger}B_{\mathrm{c}}(\Theta,0)P\right].$ (S8)
Here, we use $B_{\mathrm{c}}$ to represent the exponential of the quadratic coefficient matrix:
$B_{\mathrm{c}}(\tau_{2},\tau_{1})\equiv\prod^{\tau_{2}/\Delta\tau}_{m=\tau_{1}/\Delta\tau}~{}\mathrm{e}^{T}~{}\mathrm{e}^{V_{\mathrm{c}(m)}}.$ (S9)
The expression on the right side of Eq. (S8) has integrated out the fermion operators, replacing the Grassmann numbers and fermion statistics with a determinant representation which is computationally tractable. All matrix operations can be performed directly on a computer.
Ultimately, our Monte Carlo sampling is conducted over space-time configurations. Numerically, the weight of a space-time configuration is $A_{\mathrm{c}}~{}\mathrm{det}\left[P^{\dagger}B_{\mathrm{c}}(\Theta,0)P\right]$. Following the classical Markov importance sampling method, we continuously make tentative flips to the local components of this $d+1$ dimensional auxiliary field. We then employ the Metropolis algorithm to calculate the probability of accepting these changes based on the ratio of configuration weights before and after the flip. Specifically, we need to compute the following weight ratio:
$R_{\mathrm{c^{\prime}c}}\equiv\frac{A_{\mathrm{c^{\prime}}}~{}\mathrm{det}\left[P^{\dagger}B_{\mathrm{c^{\prime}}}(\Theta,0)P\right]}{A_{\mathrm{c}}~{}\mathrm{det}\left[P^{\dagger}B_{\mathrm{c}}(\Theta,0)P\right]},$ (S10)
where c represents the flipped configuration and c represents the original configuration. In fact, we do not need to compute the weights of the two configurations separately. This is because the flipping we perform is localized in space-time, so
$B_{\mathrm{c^{\prime}}}(\Theta,0)=B_{\mathrm{c}}(\Theta,\zeta)\left(\mathbf{1}+\Delta_{\mathrm{c^{\prime}c}}\right)B_{\mathrm{c}}(\zeta,0).$ (S11)
Here, $\Delta_{\mathrm{c^{\prime}c}}$ is a highly sparse matrix, where only the matrix elements corresponding to the degrees of freedom involved in the local auxiliary field flipping are non-zero. Thus, the ratio of the two determinants can be expressed as:
$\frac{\mathrm{det}\left[P^{\dagger}B_{\mathrm{c^{\prime}}}(\Theta,0)P\right]}{\mathrm{det}\left[P^{\dagger}B_{\mathrm{c}}(\Theta,0)P\right]}=\mathrm{det}\left\{\mathbf{1}+\Delta_{\mathrm{c^{\prime}c}}B_{\mathrm{c}}(\zeta,0)P\left[P^{\dagger}B_{\mathrm{c}}(\Theta,0)P\right]^{-1}P^{\dagger}B_{\mathrm{c}}(\Theta,\zeta)\right\}.$ (S12)
Due to the sparsity of $\Delta_{\mathrm{c^{\prime}c}}$, the determinant on the right side of the above equation only requires consideration of a few degrees of freedom involved in the flipping during computations.
In DQMC, to compute the physical observables, we only need to statistically analyze the configurational observable $\braket{O(\tau)}_{\mathrm{c}}$.
$\braket{O(\tau)}=\sum_{\mathrm{c}}\mathrm{Pr_{c}}\braket{O(\tau)}_{\mathrm{c}}+\mathcal{O}\left(\Delta\tau^{2}\right),$ (S13)
where $\mathrm{Pr_{c}}$ represents the configuration probability,
$\mathrm{Pr_{c}}=\frac{1}{Z}A_{\mathrm{c}}~{}\mathrm{det}\left[P^{\dagger}B_{\mathrm{c}}(2\tau+\tau_{0},0)P\right],$ (S14)
$\braket{O(\tau)}_{\mathrm{c}}=\frac{\bra{\psi_{0}}U_{\mathrm{c}}(2\tau,\tau)~{}O~{}U_{\mathrm{c}}(\tau,0)\ket{\psi_{0}}}{\bra{\psi_{0}}U_{\mathrm{c}}(2\tau,0)\ket{\psi_{0}}}=\frac{\bra{\psi_{T}}\mathrm{e}^{-\tau_{0}H_{\mathrm{initial},\mathrm{c}}}U_{\mathrm{c}}(2\tau,\tau)~{}O~{}U_{\mathrm{c}}(\tau,0)\mathrm{e}^{-\tau_{0}H_{\mathrm{initial},\mathrm{c}}}\ket{\psi_{T}}}{\bra{\psi_{T}}\mathrm{e}^{-\tau_{0}H_{\mathrm{initial},\mathrm{c}}}U_{\mathrm{c}}(2\tau,0)\mathrm{e}^{-\tau_{0}H_{\mathrm{initial},\mathrm{c}}}\ket{\psi_{T}}}.$ (S15)
Since we employ importance sampling, the sampling frequency is proportional to the configuration probability. Ultimately, when calculating the observable, we simply take the average over the sampled configurational observables. If the observable is a single-particle operator, meaning it can be expressed as a quadratic form of fermion operators, then one can integrate out the fermion degrees of freedom in a manner similar to Eq. (S8) and numerically compute using determinants. For observables of four-fermion operators or higher, we compute using the fermion equal-time Green’s function based on Wick’s theorem. After numerically integrating out the fermion degrees of freedom, the fermion equal-time Green’s function can be expressed using the following matrix element:
$\braket{c_{x_{1}}^{\dagger}c_{x_{2}}}_{\mathrm{c}}=\left\{B_{\mathrm{c}}\left(\frac{\Theta}{2},0\right)P\left[P^{\dagger}B_{\mathrm{c}}(\Theta,0)P\right]^{-1}P^{\dagger}B_{\mathrm{c}}\left(\Theta,\frac{\Theta}{2}\right)\right\}_{x_{1},x_{2}}.$ (S16)
@@ III RESULTS OF DIMENSIONLESS QUANTITY
where $\Delta\bm{q}$ is the minimum lattice momentum and $S({\bm{q}})$ is the structure factor.
Figure S1: Log-log plots of $R_{S}$ versus $R$ for different $L$ driven to $U_{c}=3.85$, for model (1). (a1) displays the results driven from the DSM phase. (a2) shows rescaling is applied to the horizontal axis of (a1). (b1) displays the results driven from the AFM phase. (b2) shows rescaling is applied to the horizontal axis of (b1).
Figure S2: Log-log plots of $R_{S}$ versus $R$ for different $L$ driven to $V_{c}=1.355$, for model (4). (a1) displays the results driven from the DSM phase. (a2) shows rescaling is applied to the horizontal axis of (a1). (b1) displays the results driven from the CDW phase. (b2) shows rescaling is applied to the horizontal axis of (b1).
@@
where $S_{i}^{z}$ is the staggered magnetization operator defined as $S_{i}^{z}\equiv\vec{c}^{\dagger}_{i,A}\sigma^{z}\vec{c}_{i,A}-\vec{c}^{\dagger}_{i,B}\sigma^{z}\vec{c}_{i,B}$ with $i$ characterizes cells and ${\vec{c}^{\dagger}\equiv({c_{\uparrow}^{\dagger},c_{\downarrow}^{\dagger}}})$.

### v2

@@ I IMAGINARY-TIME DRIVEN DYNAMICS
In the main text, we explored the driven dynamics of the Hubbard model (1) and the $t$-$V$ model (4). Here, we elucidate the theory of imaginary-time driven dynamics.
For model (1) and (4), we linearly vary the interaction strength $U~{}(V)$ with imaginary time variable $\tau$ at the rate $R$ as
$U=U_{0}\pm R\tau,$ (S1)
$V=V_{0}\pm R\tau,$ (S2)
where the +()+(-)+ ( - ) represents driving from disordered (ordered) initial state at $U_{0}~{}(V_{0})$. In our numerical simulation, for disordered initial state, we set $U_{0}=0,~{}V_{0}=0$, while for ordered initial state we set $U_{0}=11.85,~{}V_{0}=2.5$ far from the critical point.
The wave function $|\psi(\tau)\rangle$ obeys the imaginary-time Schrödinger equation De Grandi et al. (2011); Grandi et al. (2013); Liu et al. (2013)
@@
The formal solution of Eq. (S3) is given by $|\psi(\tau)\rangle=U(\tau,0)|\psi(0)\rangle$ in which time evolution operator
[…]
@@ II Determinant quantum Monte Carlo
We implement DQMC simulation Assaad and Evertz (2008) to explore the imaginary-time driven dynamics of our models. Starting with a trial wave function |$\ket{\psi_{T}}$, we project the system onto the desired initial state |$\ket{\psi(0)}$ with a sufficiently long projection time $\tau_{0}$. That is $|\psi(0)\rangle=\mathrm{e}^{-\tau_{0}H_{0}}|\psi_{T}\rangle$ with $H_{0}$ the Hamiltonian of $|\psi(0)\rangle$ corresponding to $V_{0}$. Then we follow Eq. (S2) to vary $V$ and observe the scaling behavior of observables at final state, at which the expectation value of observables is given by
$\braket{O(\tau)}=\frac{\bra{\psi(\tau)}O\ket{\psi(\tau)}}{\left\langle\psi(\tau)\middle|\psi(\tau)\right\rangle\ }.$ (S5)
We write the Hamiltonian of model (1) and model (4) as
$H=H_{t}+H_{V},$ (S6)
where $H_{t}$ is the kinetic energy and $H_{V}$ is the interaction term. In DQMC, one usually employs Trotter decomposition to discretize imaginary time and discrete Hubbard-Stratonovich (HS) transformation to split the two-body fermion-fermion coupling term into a one-body fermion-auxiliary field coupling.
The Trotter decomposition in driven dynamics is expressed as
[…]
where the imaginary-time slice $\Delta\tau=\tau/N_{\tau}$ and $\tau_{n}=n\Delta\tau$.
We apply the following HS transformation to model (1):
@@
$\mathrm{e}^{-\frac{\Delta\tau U(\tau_{n})}{2}\left(n_{i\uparrow}+n_{i\downarrow}-1\right)^{2}}=\sum_{l_{i,\tau_{n}}=\pm 1,\pm 2}\gamma(l_{i,\tau_{n}})\mathrm{e}^{\mathrm{i}\sqrt{\frac{\Delta\tau U(\tau_{n})}{2}}\eta(l_{i,\tau_{n}})\left(n_{i\uparrow}+n_{i\downarrow}-1\right)},$ (S8)
and for model (4):
$\mathrm{e}^{-\frac{\Delta\tau V(\tau_{n})}{2}\left(n_{i}+n_{j}-1\right)^{2}}=\sum_{l_{i,\tau_{n}}=\pm 1,\pm 2}\gamma(l_{i,\tau_{n}})\mathrm{e}^{\mathrm{i}\sqrt{\frac{\Delta\tau V(\tau_{n})}{2}}\eta(l_{i,\tau_{n}})\left(n_{i}+n_{j}-1\right)},$ (S9)
where the four-component space-time auxiliary fields $\gamma$ and $\eta$ take the following values:
$\gamma(\pm 1)=1+\sqrt{6}/3,\ \gamma(\pm 2)=1-\sqrt{6}/3,$ (S10)
$\eta(\pm 1)=\pm\sqrt{2\left(3-\sqrt{6}\right)},\ \eta(\pm 2)=\pm\sqrt{2\left(3+\sqrt{6}\right)}.$ (S11)
After the aforementioned operations, the imaginary-time evolution operator can be expressed in quadratic form of fermion operators:
$U(\tau,0)=\sum_{l}A_{l,\tau_{n}}\prod^{N_{\tau}}_{n=1}~{}\mathrm{e}^{\vec{c}^{\dagger}K\vec{c}}~{}\mathrm{e}^{\vec{c}^{\dagger}W_{l,\tau_{n}}\vec{c}},$ (S12)
where $K$ and $W_{l,\tau_{n}}$ are the quadratic coefficient matrices of hopping term and interaction term respectively, and $A_{l,\tau_{n}}$ is the coefficient. Then the fermionic degrees of freedom can be integrated out and we obtain
$\bra{\psi(0)}U^{\dagger}(\tau,0)U(\tau,0)\ket{\psi(0)}=\sum_{l}A_{l,\tau_{n}}~{}\mathrm{det}\left[P^{\dagger}B_{l}(2\tau,0)P\right],$ (S13)
where $P$ represents the initial state with fermions operators intergrated out and
$B_{l}(2\tau,0)\equiv\prod^{2N_{\tau}}_{n=1}~{}\mathrm{e}^{K}~{}\mathrm{e}^{W_{l,\tau_{n}}}.$ (S14)
Now we can perform matrix operations directly on a computer and sample over the space-time configurations to evaluate the expectation value of physical observables. The subsequent DQMC process is the same as that for equilibrium state approach Assaad and Evertz (2008).
@@ III RESULTS OF DIMENSIONLESS QUANTITY
where $S({\bm{q}})$ is the structure factor and $\Delta\bm{q}$ is the minimal lattice momentum.
Figure S1: Log-log plots of $R_{S}$ versus $R$ for different $L$ driven to $U_{c}=3.85$, for model (1). (a) displays the results driven from the DSM phase before (a1) and after (a2) rescaling. (b) displays the results driven from the AFM phase before (b1) and after (b2) rescaling.
Figure S2: Log-log plots of $R_{S}$ versus $R$ for different $L$ driven to $V_{c}=1.355$, for model (4). (a) displays the results driven from the DSM phase before (a1) and after (a2) rescaling. (b) displays the results driven from the CDW phase before (b1) and after (b2) rescaling.
@@
where the staggered magnetization operator $S_{i}^{z}\equiv\vec{c}^{\dagger}_{i,A}\sigma^{z}\vec{c}_{i,A}-\vec{c}^{\dagger}_{i,B}\sigma^{z}\vec{c}_{i,B}$ with $i$ characterizes cells and ${\vec{c}^{\dagger}\equiv({c_{\uparrow}^{\dagger},c_{\downarrow}^{\dagger}}})$.